# Per-example gradients: a new frontier for understanding and improving optimizers

### v1

@@ Abstract
Training algorithms in deep learning usually treat a mini-batch of samples as a single object; they average gradients over the mini-batch, and then process the average in various ways. Computing other statistics beyond the average may have been seen as prohibitively resource intensive in automatic differentiation (AD) frameworks. We show that this is not the case. Generally, gradient statistics can be implemented through a surgery of the AD graph, which, in some cases, incur almost no computational and memory overheads compared to the mini-batch gradient computation. Additionally, we show that in certain classes of models, including transformers, JAX’s vectorization transformation offers a viable implementation for prototyping and experimentation. We then revise our understanding of two nonlinear operations in optimization through the lens of per-example gradient transformations. We first study signSGD and show that the optimal placement of the sign operation in the gradient processing chain is crucial to success and can be predicted with a simple signal-to-noise ratio argument. Next we study per-example variations of the Adam preconditioner, and show that optimization is best served when the preconditioner is dominated by the mean rather than the variance of the gradient distribution — in contrast to conventional wisdom. Overall we demonstrate that per-example gradient information enables new analyses and possibilities for algorithm design.
@@ 1 Introduction
The success of modern machine learning would not be possible without efficient methods of computing gradients of loss functions through neural networks, primarily through the invention of reverse-mode automatic differentiation (AD) (Linnainmaa, 1970). With the advent of accelerators like GPUs, practitioners are now able to approximate the expected gradients of a model’s loss function on a dataset by averaging gradients on large batches of data. This has allowed model and data complexity to grow rapidly while maintaining tractability of training algorithms (Hoffmann et al., 2022).
A key feature of reverse-mode AD is that it never stores the gradients of individual elements of a batch. This increases the memory efficiency of the process, but makes it impossible to answer questions about the distribution of per-example gradients, or allow for optimizers to depend on higher order moments of the gradient distribution. This leaves a vast part of training algorithm design space inaccessible for researchers, particularly if they want to work at large scales.
Access to per-example gradient covariance statistics is useful for understanding deep learning; this information can be used to predict quantitative properties of loss trajectories at scale (Yin et al., 2018; Qiu et al., 2025). There is practical interest in per-example gradient transformations as well. There is some evidence that optimizers which have access to per-example gradients can have better stability and predictability than typical optimizers (Wang and Aitchison, 2024). The growing interest for distributed optimizers like DiLoCo (Douillard et al., 2023) calls for studying how optimizers may benefit from aggregating more than just the gradients (Zhang et al., 2023). This has led to attempts to efficiently compute per-example gradient norms (Kong and Munoz Medina, 2023; Bu et al., 2023), to come up with ways to modify AD more generally to compute higher order moments of gradient distributions (Dangel et al., 2019), or to develop AD frameworks capable of leveraging broadcast and reduce sum operations at any scale (Rush et al., 2024).
Motivated by these examples, we study the technical challenges arising from computing generic gradient statistics. We show the following:
  • Generic gradient statistics can be implemented with a computational graph surgery, and, in some cases, gradients’ structure allows the approach to be as efficient as computing average gradients (Section 2.2).
  • For sequence-based models, an implementation using just-in-time compiled JAX vmap vectorization transformation only incurs a moderate computational overhead, providing a path for prototyping (Section 2.3).
Through these technical findings we do not claim that generic gradient statistics are inexpensive, we simply claim that they are not prohitively expensive. We can then use gradient statistics to improve our understanding of two optimization algorithms:
@@
  • Adam (Section 3.2). We study Adam variants which operate on per-example statistics like the second moment. We provide, to the best of our knowledge, the first true implementation of these algorithms at scale, have the predicted scaling properties with batch size. We show that preconditioners which depend on the mean squared train faster and more stably than those which depend on the variance, in contrast to conventional wisdom.
@@ 2.1 Objective and challenges
    - $\texttt{fetch\_batch(train\_data)}=x_{1},\ldots,x_{B}$
@@
    - $\texttt{loss(params, batch)}=\frac{1}{B}\sum_{i=1}^{B}f(\theta;x_{i})$
@@
    - $\texttt{grad(loss)(params, batch)}=\nabla\frac{1}{B}\sum_{i=1}^{B}f(\theta;x_{i})=\frac{1}{B}\sum_{i=1}^{B}\nabla f(\theta;x_{i})$
@@
where $x_{1},\ldots,x_{B}$ are assumed to be i.i.d. samples of $X$. However, other statistics might be of interest for understanding existing algorithms, and to design new ones. We may want to approximate the expectation of an arbitrary function $\phi$ of the gradients in a minibatch as
@@
Specific examples include the elementwise square $\mathbb{E}[\nabla f(\theta;X)^{2}]$, the elementwise variance $\mathbb{E}[(\nabla f(\theta;X)-\mathbb{E}[\nabla f(\theta;X)])^{2}]$, and sign $\mathbb{E}[{\rm sign}(\nabla f(\theta;X))]$.
If $\frac{1}{B}\sum_{i=1}^{B}\phi(\nabla f(\theta;x_{i}))$ is the gradient of some function f, it could be computed by automatic differentiation as grad(f)(params, batch). However, this is generally not the case. With basic access to gradient oracle calls like grad(loss)(params, sample), there are two simple ways to compute this statistic, both at the ends of a computation-memory trade-off.
@@
    Memory intensive implementation. Compute $\nabla f(\theta;x_{1}),\ldots,\nabla f(\theta;x_{B})$ in parallel, apply $\phi$ to each in parallel, and then take the average. Assuming that the statistics we search are of lower or equal dimension than the gradients, i.e., $\dim(\phi(\nabla f(\theta;x_{i})))\leq\dim(\nabla f(\theta;x_{i}))$, this requires $BP$ memory storage, with $P$ the dimension of the parameters, and can be prohibitive for some architectures and hardware.
@@
    Computationally intensive implementation. Accumulate the average $\bar{\phi}$ in a loop for $i=1$ to $B$, computing the gradient, $\nabla f(\theta;x_{i})$, applying $\phi(\nabla f(\theta;x_{i}))$, and updating $\bar{\phi}\leftarrow[(i-1)\bar{\phi}+\phi(\nabla f(\theta;x_{i}))]/i$. Assuming again $\dim(\phi(\nabla f(\theta;x_{i})))\leq\dim(\nabla f(\theta;x_{i}))$, this prevents using more memory storage than the one needed to compute a gradient. However, it requires $B$ more calls to a gradient oracle.
Efficiently implementing these non-linear averages requires a dive into the mechanics of AD, and the computational bottlenecks of deep learning architectures.
@@
2.2 Gradient statistics by computational graph surgery
In the computation of the mini-batch gradient, most computations preserve per-example information, as can be seen in the generic computational graph (Figure 1). The averaging of the gradients over the batch is generally the last operation (illustrated in Figure 1, detailed in Appendix B.1). If we parse the computational graph of the minibatch gradient, we can “inject” the desired computation $\phi$ to the individual gradients before averaging. In the remainder of this section we describe an efficient way to perform this operation.
For certain operations $\phi$ can be injected with negligible overhead. Specifically, consider a dense layer parameterized by weights $W_{k}\in\mathbb{R}^{D\times D}$. In the computation of the loss $f(\theta;x_{i})$, this layer acts on intermediate inputs $s_{i}$ as $t_{i}^{\top}=s_{i}^{\top}W_{k}$. To compute the gradient of $f(\theta,x_{i})$ w.r.t. $W_{k}$, reverse mode AD performs a backward pass through the transpose Jacobians that computes $r_{i}\coloneqq\nabla_{t_{i}}f(\theta;x_{i}),$ the derivative of $f(\theta;x_{i})$ w.r.t. the output of that layer. The gradient w.r.t. $W_{k}$ is then computed by passing $r_{k}$ through the adjoint operation of $W_{k}\mapsto s_{i}^{\top}W_{k}$. This gives $\nabla_{W_{k}}f(\theta;x_{i})=s_{i}r_{i}^{\top}$ – a rank-one matrix – for a single example. For the mini-batch gradient we get
$\nabla_{W_{k}}\sum_{i=1}^{B}f(\theta;x_{i})=\sum_{i=1}^{B}s_{i}r_{i}^{\top}=SR,$ (3)
where $S=(s_{1},\ldots,s_{B})^{\top}\in\mathbb{R}^{B\times D}$ and $R=(r_{1},\ldots,r_{B})\in\mathbb{R}^{D\times B}$, stack the intermediate inputs and intermediate derivatives computed in the independent computational paths illustrated in Figure 1. The required memory of this operation is $O(BD+D^{2})$.
The rank-one nature of the individual gradients can be exploited by any “factorable” operation $\phi$ where $\phi(sr)=\phi(s)\phi(r)$. Examples of such functions include the sign  function as well as any power $\phi(s)=s^{\alpha}$. Of particular interest is the elementwise square of the gradient which amounts to
$\sum_{i=1}^{B}[\nabla_{W_{k}}f(\theta;x_{i})]^{2}=\sum_{i=1}^{B}(s_{i}r_{i}^{\top})^{2}=\sum_{i=1}^{B}(s_{i})^{2}(r_{i})^{2}{}^{\top}=S^{2}R^{2}.$ (4)
@@ 2.2 Gradient statistics by computational graph surgery
Apart from an elementwise square of the matrices $S,R$, this operation has the same memory and computational complexity as computing the sum of gradients. It can take advantage of accelerators, and it does not suffer from the $O(BD^{2})$ memory cost of the naive memory intensive implementation from Section 2.1.
@@
The overall approach amounts to a computational graph surgery and can be implemented in JAX (Bradbury et al., 2018); we provide code and benchmarking in Appendix B. For each operation of interest $\phi$, the method modifies the basic linear primitives in JAX. We note that a related approach was taken on a different problem (fast computation of NTK) in Novak et al. (2022). The specific case of efficient computation of squared gradients has been pointed out previously by Dangel et al. (2019), but their approach required overriding derived classes (specifically, base layers) in PyTorch (Paszke et al., 2019). In contrast our approach modifies basic framework operations, allowing automatic compatibility with downstream modules.
Figure 2: Memory footprint along program execution as reported by a code profiler of the train step of the usual Adam algorithm and its per-example variant, MicroAdam (Section 3.2), for a 1.2B transformer in Nanodo (Liu et al., 2024). The peak memory corresponds to the accumulation of the memory during the forward pass of automatic differentiation necessary to compute gradients. We observe that the per-example variant may incur more operations (longer tail) that translate in longer execution time. But the peak memory is the same.
2.3 Opportunities in sequence-level architectures
While the computational graph surgery approach is demonstrably efficient, it may not be needed for tractable prototyping and exploration in certain settings. In particular, more naive implementations are viable in sequence to sequence models like transformers.
For dense layers taking vector inputs, processing a batch of data of size $B$ with a layer of width $D$ contributes to the memory cost of the backwards pass in two ways: a memory cost of $BD$ from activations, and a cost of $D^{2}$ from the parameters. Historically, the parameter cost $D^{2}$ was often limiting, which meant that storing the individual gradients at a cost $BD^{2}$ was prohibitively expensive.
However, for sequence-based layers trained on sequences represented by $L$ vectors of size $D$, the cost of storing the activations of the batch alone is $BDL$, while the parameters take $D^{2}$ space. If $L\geq D$, then the cost of storing $B$ individual gradients is less than the memory cost of the activations! This does not apply to the embedding and final head layers but it still suggests that gradient statistics may be computable in these models. So to test the hypothesis, we implemented the elementwise square of the gradient of a transformer model in JAX using the automatic vectorization tool vmap to implement the per-example gradients statistics. We found that computing these statistics did not increase the peak memory usage (which was reached by checkpointing activations, see Figure 2) and it only added a small compute overhead ($\approx 17\%$) on a model of size 1.2B parameters with TPU v5. Here JAX may also benefit highly from the underlying XLA’s optimizations (Appendix A) that can best trade-off computational and memory costs.
Overall, our results suggest that for transformer architectures and some other layers detailed in Appendix B.4, the vmap approach offers a reasonably efficient way to access non-linear gradient statistics and prototype new optimizers or diagnostic tools. The computational graph surgery can complete the prototyping to generalize these optimizers to other architectures while keeping a small computational overhead.
3 New perspectives using mini-batch gradient statistics

### v2

@@ Abstract
When computing gradients, deep learning training algorithms typically treat the mini-batch as a fundamental unit — only returning batch-averaged gradients. Computing non-linear statistics of the mini-batch gradient distribution has traditionally been viewed as prohibitively expensive or requiring complex, custom implementations. We challenge this view by demonstrating that sequence-level architectures offer a natural testbed for prototyping algorithms based on per-example gradients. We show that staged programming languages like JAX enable generic manipulations of mini-batch gradient computations. We then build on dangel2019backpack to derive implementations of specific per-example or per-token operations with negligible computational or memory overhead. Finally, we leverage our findings to re-examine two nonlinear optimization operations. First, we analyze signSGD, showing that the optimal placement of the sign operation is critical to success and can be predicted via a simple signal-to-noise ratio argument. Second, we investigate per-example variations of the Adam preconditioner and find that, contrary to conventional wisdom, optimization is best served when the preconditioner is dominated by the mean squared of the gradient distribution rather than its variance. Overall our work shows that accessible per-example gradient information unlocks new avenues for algorithm analysis and design.
Machine Learning, ICML
@@ 1 Introduction
The success of modern machine learning was enabled by efficient methods of computing gradients of loss functions through neural networks, primarily through the invention of reverse-mode automatic differentiation (AD) (linnainmaa1970representation). With the advent of accelerators like GPUs, practitioners are now able to approximate the expected gradients of a model’s loss function on a dataset by averaging gradients on large batches of data. This has allowed model and data complexity to grow rapidly while maintaining tractability of training algorithms (hoffmann2022training).
In recent years, there has been growing interest in accessing and using more complex information about the gradient distribution. For example, per-example gradient covariance statistics is useful for understanding deep learning; this information can be used to predict quantitative properties of loss trajectories at scale (yin2018gradient; mccandlish2018empirical; liu2020understanding; faghri2020study; gray2024normalization; qiu2025scaling). There is practical interest in per-example gradient transformations as well. There is some evidence that optimizers which have access to per-example gradients can have better stability and predictability than typical optimizers (wang2024batch). The growing interest in distributed optimizers like DiLoCo (douillard2023diloco) calls for study of optimizers that operate beyond just batch-averaged gradients (zhang2023adam).
However, a key feature of reverse-mode AD is that it never stores the gradients of individual elements of a batch. This increases the memory efficiency of the process, but makes it impossible to answer questions about the distribution of per-example gradients, or allow for optimizers to depend on higher order moments of the gradient distribution. This leaves a vast part of training algorithm design space inaccessible for researchers, especially at large scales. There have been attempts to efficiently compute statistics per-example gradient norms (kong2023unified; bu2023differentially), and to come up with ways to modify AD more generally to compute higher order moments of gradient distributions (dangel2019backpack; schneider2021cockpit), but general development remains difficult.
Motivated by these examples, we study the technical challenges arising from computing generic gradient statistics. We show the following (Section 2):
  • In some workloads, including transformers, simple automatic vectorization tools like vmap in JAX (jax2018) allow for quick prototyping without incurring prohibitive overheads,
  • Per-example or “per-token” gradient statistics may be implemented with a negligible memory and computational overhead for promising algorithms and metrics on specific architecture families.
Our results suggest that computing generic gradient statistics is not prohibitively expensive, and in some cases come at virtually no overhead. We then use gradient statistics to improve our understanding of two optimization algorithms:
@@
  • Adam (Section 3.2). We study Adam variants which operate on per-example statistics like the second moment. We provide, to the best of our knowledge, the first true implementation of these algorithms at scale, confirming the predicted scaling properties with batch size. We show that preconditioners which depend on the mean squared train faster and more stably than those which depend on the variance, in contrast to conventional wisdom.
@@ 2.1 Objective and challenges
    $\small{\texttt{fetch\_batch(train\_data)}=x_{1},\ldots,x_{B}}$
@@
    $\small{\texttt{loss(params, batch)}=\frac{1}{B}\sum_{i=1}^{B}f(\theta;x_{i})}$
@@
    $\small{\texttt{grad(loss)(params, batch)}=\nabla\frac{1}{B}\sum_{i=1}^{B}f(\theta;x_{i})}$
@@
where $x_{1},\ldots,x_{B}$ are assumed to be i.i.d. samples of $X$.
However, other statistics might be of interest for metrics, scientific exploration, and new algorithm design. Consider the problem of estimating the expectation of an arbitrary function $\phi$ of the gradient distribution using the mini-batch:
@@
Statistics of interest include the element-wise variance $\mathbb{E}[(\nabla f(\theta;X)-\mathbb{E}[\nabla f(\theta;X)])^{2}]$, the sign $\mathbb{E}[{\rm sign}(\nabla f(\theta;X))]$, or clipping $\mathbb{E}[\mathrm{clip}(\nabla f(\theta;X))]$.
As a running example, we will consider MicroAdam— a variant of Adam which uses the average element-wise square gradients over a batch rather than the squared average gradients for the preconditioner (wang2024batch). Formally, denoting per-example gradients $g_{i}=\nabla f(\theta,x_{i})$, the preconditioner becomes
$\nu_{\rm micro}\coloneqq\frac{1}{B}\sum_{i=1}^{B}g_{i}^{2},\ \text{in lieu of}\ \nu_{\rm adam}\coloneqq\left(\frac{1}{B}\sum_{i=1}^{B}g_{i}\right)^{2}.$ (3)
The rest of the update rule remains the same. We will benchmark this algorithm against standard Adam to evaluate our methods, and will analyze the algorithm in detail in Section 3.2.
With basic access to gradient oracle calls like grad(loss)(params, sample), there are two simple ways to compute $\nu_{\text{micro}}$, both at the ends of a computation-memory trade-off.
@@
    Memory intensive implementation. Compute and square $\nabla f(\theta;x_{1}),\ldots,\nabla f(\theta;x_{B})$ in parallel, then take the average. This requires $BP$ memory storage, with $P$ the dimension of the parameters, and can be prohibitive for some architectures and hardware.
@@
    Computationally intensive implementation. Accumulate the average $\nu_{\rm micro}$ in a loop for $i=1$ to $B$, computing the gradient, $\nabla f(\theta;x_{i})$, square it, and update $\nu_{\rm micro}\leftarrow[(i-1)\nu_{\rm micro}+\nabla f(\theta;x_{i})^{2}]/i$. This prevents using more memory storage than the one needed to compute two gradients. However, it requires $B$ more calls to a gradient oracle.
In the remainder of the section, we will show that there are opportunities to improve on these naive methods by detailed examination of the compute flow during AD.
2.2 Memory bottlenecks in deep learning
The memory required to compute statistics of the form (2) does not reside only in the final $B$ gradients $\nabla f(\theta;x_{i})$. Automatic differentiation (AD) first checkpoints intermediate values in a forward pass. This leads to a simple fact that underpins potential opportunities to easily implement nonlinear statistics like (2):
Fact 2.1.
For layers where the input size exceeds the parameter size, the memory typically reserved for input checkpoints can be repurposed to temporarily store $B$ individual gradients. This a priori enables the computation of nonlinear gradient statistics (2), such as $\nu_{\rm micro}$, without increasing peak memory usage.
To better understand this, we review some standard layers using Einstein summation (einsum) notation:
  1. 1.
    Dense layers in standard MLPs ($\mathbf{D,DF\rightarrow F}$): Here, the parameter size, $DF$ is larger than the input size $D$. Storing individual gradients at a cost $BDF$ is larger than the cost of checkpointing inputs, $BD$, and is generally prohibitively expensive.
  2. 2.
    Dense layers in transformers ($\mathbf{LD,DF\rightarrow LF}$): Here as soon as the length $L$ is larger than the hidden dimension $F$, the cost of storing $B$ individual gradients is actually less than the memory cost of the activations!
Figure 1: Memory footprint along program execution as reported by a code profiler of a train step. We compare the usual Adam algorithm and its per-example variant, MicroAdam (Section 3.2) implemented in JAX using the automatic vectorization tool vmap. The peak memory corresponds to the accumulation of the memory during the forward pass of automatic differentiation necessary to compute gradients. We observe that the per-example variant may incur more operations (longer tail) that translate in longer execution time. But the peak memory is the same.
Figure 1 shows the cost in memory incurred for Adam or its per-example variant using $\nu_{\rm micro}$, MicroAdam, on a 1.2B transformer in Nanodo (nanodo) using TPU v5e. It depicts a usual staircase shape arising from checkpointing (blondel2024elements, Figure 8.4). It confirms that, for a transformer, the peak memory remains almost unchanged. It also hints that the first layer could also benefit from all freed memory of later layers. Fact 2.1 only points out potential room for individual layers; the total cost depends on the global architecture and compilation optimizations that can trade-off memory and speed. A convolutional network may for example have some per-example-friendly layers like convolutions but also usual dense layers for which the cost is not negligible.
Our analysis suggests that modern just-in-time compilers may be able to automatically identify and exploit this structure in some architectures. In order to test this hypothesis, we implemented MicroAdam across $7$ workloads of the AlgoPerf challenge (kasimbeg2025accelerating) using the JAX library’s vmap vectorization capability, see dahl2023benchmarking for the workloads details. We compared the peak memory and time per step to Adam (Figure 2). In the transformer model workloads for text or vision (first three), the increase in time or memory cost of the MicroAdam algorithm is modest, especially compared to the per-device batch size (see Appendix E.1). It illustrates that experimentation with per-example methods on transformer is not prohibitively expensive. On the other workloads we do not observe excessive increase in memory used (at most $2\times$) but this modest increase in memory may be paid by a non-negligible increase in computational cost as compilers may be able to automatically optimize for memory constraints with adequate fusions of some operations
Figure 2: Relative efficiency of MicroAdam compared to Adam in terms of time and memory costs across the workloads of the AlgoPerf Challenge (dahl2023benchmarking). On the three first workloads (transformers) time and memory costs overhead are modest and allow for quick prototyping of per-example methods. On the other workloads, memory requirements remain bounded to 2x the cost for Adam but may trade this limited increase with substantial time increase.
@@ 2.2 Memory bottlenecks in deep learning
2.3 Computational graph surgery
While vmap already provides us with a useful method across multiple workloads, we can improve even further by diving into the computational graphs generated by AD. Per-example information is preserved across most of the operations in the computation graph; the averaging of the gradients over the batch is generally the last operation (illustrated in Figure 3, detailed in Appendix B.1). If we parse the computational graph of the minibatch gradient, we can “inject” the desired computation $\phi$ to the individual gradients before averaging. For some operations $\phi$ like computing $\nu_{\rm micro}$, and some layers reviewed below the injection can be done with negligible overhead.
Dense layers in standard MLPs.
Take a dense layer in a standard MLP, parameterized by $W\in\mathbb{R}^{D\times F}$, which transforms a mini-batch of incoming activations $H\in\mathbb{R}^{B\times D}$ to outputs $Z\in\mathbb{R}^{B\times F}$. To compute the mini-batch gradient $\sum_{b}\nabla_{W}f(\theta;x_{b})\eqqcolon G\in\mathbb{R}^{D\times F}$ of the loss w.r.t. $W$, reverse mode AD performs a backward pass through transpose Jacobians across the network that computes the derivatives $\Lambda\in\mathbb{R}^{B\times F}$ of the loss w.r.t. the output of that layer. In einsum notations, the forward computation defining this layer and the computation leading to the gradient are
$\begin{array}[]{cccc}\mathrm{BF}&\leftarrow&\mathrm{BD},&\mathrm{DF}\\ Z&=&H&W,\end{array}\qquad\begin{array}[]{cccc}\mathrm{DF}&\leftarrow&\mathrm{BD},&\mathrm{BF}\\ G&=&H&\Lambda,\end{array}$
Both incoming activations $H$ and co-tangents $\Lambda$ are computed in independent computational paths illustrated in Figure 3. The mini-batch gradient $G$ agglomerates these paths in a sum of rank-one vectors $G=\sum_{b}h_{b}\lambda_{b}^{\top}$.
Computing the average elementwise squared gradients, can be done with a simple modification as we have $\sum_{b}\nabla_{W}f(\theta;x_{b})^{2}=\sum_{b}(h_{b}\lambda_{b}^{\top})^{2}=\sum_{b}h_{b}^{2}\lambda_{b}^{2}{}^{\top}$, i.e.,
$\begin{array}[]{ccccc}&\mathrm{DF}&\leftarrow&\mathrm{BD},&\mathrm{BF}\\ \sum_{b}\nabla f(\theta;x_{b})^{2}\eqqcolon&S&=&H^{\odot 2}&\Lambda^{\odot 2},\end{array}$
@@ 2.3 Computational graph surgery
where $M^{\odot 2}$ denotes the elementwise square of a tensor $M$. Apart from squaring $H,\Lambda$, the result has the same memory cost, $O(BD+D^{2})$, and computational complexity as computing the sum of gradients. It can take advantage of accelerators, and does not suffer from the $O(BD^{2})$ memory cost of a naive memory-intensive implementation ( Section 2.1).
This computation of second moments for MLPs was first presented by dangel2019backpack. It can be generalized to any operation $\phi$ that is factorable, i.e., such that $\phi(ab)=\phi(a)\phi(b)$ by using the rank-one nature of the gradients in this case. Examples of such functions include the sign function as well as any power $\phi(s)=s^{\alpha}$. In particular this means that any Taylor series can be efficiently implemented with this methodology. These algebraic manipulations are also similar to the “ghost-clipping” trick presented by goodfellow2015efficient that compute squared norms as the einsum contraction $(\|\nabla_{W}f(\theta;x_{b})\|_{2}^{2})_{b=1}^{B}N_{b}\eqqcolon\leftarrow H_{b,d},\Lambda_{b,f}N_{b}$.
Dense layers in sequence-level architectures.
In sequence-level architectures like transformers, the forward computation and the computation leading to the gradient take the form
$\begin{array}[]{cccc}\mathrm{BLF}&\leftarrow&\mathrm{BLD},&\mathrm{DF}\\ Z&=&H&W,\end{array}\qquad\begin{array}[]{cccc}\mathrm{DF}&\leftarrow&\mathrm{BLD},&\mathrm{BLF}\\ G&=&H&\Lambda,\end{array}$
with $L$ denoting the length dimension of sequences fed to the model. Computing the average squared gradients amount to computing
$\begin{array}[]{cccc}\mathrm{BDF}&\leftarrow&\mathrm{BLD},&\mathrm{BLF}\\ M&=&H&\Lambda,\end{array}\qquad\begin{array}[]{ccc}\mathrm{DF}&\leftarrow&\mathrm{BDF}\\ S&=&M^{\odot 2},\end{array}$
i.e., it requires first reducing over the length dimension, creating a $BDF$ tensor of per-example gradients, that is squared and then reduced. The leading complexity of the resulting operations remains $O(BLDF)$ but it has also an intermediate memory cost $O(BDF)$.
An alternative is to consider per-example “per-token” second moments, i.e., computing
$\begin{array}[]{ccccc}\mathrm{DF}&\leftarrow&\mathrm{BLD},&\mathrm{BLF}&\\ S^{\prime}&=&H^{\odot 2}&\Lambda^{\odot 2}&=\sum_{bl}h_{bl}^{2}\lambda_{bl}^{2}{}^{\top},\end{array}$
As for dense layers in MLPs, such an operation comes at a negligible overhead. In a transformer, tokens attend to each-other in the attention layers, both in the forward and backward pass, because the transpose Jacobian of the attention block won’t be separable along tokens. So the elements $\lambda_{bl}$ of the cotangents do not correspond to derivatives of the loss at token $l$ of sample $b$. Nevertheless the derived quantity can be used as proxy to capture relevant statistics. Note that different einsum contractions may also lead to per-sample, “per-token” norms as $N_{bl}\leftarrow H_{bld}^{\odot 2}\Lambda_{bld}^{\odot 2}$.
Implementation
@@
To implement these operations we may either define some custom derivative rules for the layers of interest as previously proposed by dangel2019backpack in PyTorch (pytorch2019). Functional programming languages like JAX (jax2018) or Functorch (functorch2021) enable a more generic approach: a computational graph surgery that transforms the traced computation of a gradient. We provide code and benchmarking in Appendix B. For each operation of interest $\phi$, the method modifies the basic linear primitives in JAX. We note that a related approach was taken on a different problem (fast computation of NTK) in novak2022fast. Such approaches modify basic framework operations, allowing automatic compatibility with downstream modules.
3 New perspectives from gradient statistics